OpenStudy (anonymous):

What is the center of the circle that has a diameter whose endpoints are (6, -4) and (18, 10)? (12, 7) (12, 6) (12, 3)

OpenStudy (anonymous):

(12,3)

OpenStudy (anonymous):

All I did was graph the points, then I found the center.

OpenStudy (anonymous):

Okay, one more question.

OpenStudy (anonymous):

What is the center of the circle that has a diameter whose endpoints are (-3, -5) and (-11, -7)? (-8, -1) (-7, -6) (-4, -6)

OpenStudy (nurali):

\[\frac{ (x _{1}+x _{2}) }{ 2 },\frac{ y _{1}+y _{2} }{ 2 }\] (x1,y1)= (6, -4) (x2,y2)=(18, 10) \[\frac{ 6+18 }{ 2 },\frac{ -4+10 }{ 2 }\] \[\frac{ 24 }{ 2 },\frac{ 6 }{ 2 }\] \[12,3\]

OpenStudy (nurali):

Same formula u apply in second question

OpenStudy (anonymous):

(-7, -6), I did the same thing. I plotted the points, and then I found the center. I always double check using rise over run.